### v1

# Universal neural network potentials as descriptors: Towards scalable chemical property prediction using quantum and classical computers

Abstract
Accurate prediction of diverse chemical properties is crucial for advancing molecular design and materials discovery. Here we present a versatile approach that uses the intermediate information of a universal neural network potential as a general-purpose descriptor for chemical property prediction. Our method is based on the insight that by training a sophisticated neural network architecture for universal force fields, it learns transferable representations of atomic environments. We show that transfer learning with a graph neural network potential M3GNet achieves accuracy comparable to state-of-the-art methods for predicting the NMR chemical shifts ofH, C, N, O, and F using quantum machine learning as well as a standard classical regression model, despite the compactness of its descriptors. This work provides an efficient way to accurately predict properties, potentially accelerating the discovery of new molecules and materials.
I Introduction
As evidenced by the enumeration of 166.4B possible organic molecules containing up to 17 heavy elements, such as C, N, O, S, and halogens (excluding hydrogen), the expansion of the chemical space is astronomical with the increase in types and numbers of elements.[1, 2] This vast landscape has given rise to a multidisciplinary approaches to combining experimental and computational chemistry for the discovery of new chemical substances and materials in a wide range of fields, including material, catalysis, and drug design.[1, 2, 3, 4, 5, 6] Although quantum chemistry and first-principles calculations offer accurate descriptions of chemical substances, their high computational demands make an exhaustive exploration of the chemical space impractical.[7, 8, 9, 10, 11, 12, 6] However, machine- and deep-learning techniques are overcoming these limitations to enable a more extensive exploration.[13, 14, 15, 16, 17, 10, 18, 19, 20, 21, 4, 22, 23, 24, 25, 26, 27, 6]
With machine learning, physics-inspired descriptors that characterize the chemical space have been developed and serve as the cornerstone for building efficient and highly accurate models.[28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 21, 38, 23, 39, 40] Smooth overlap of atomic positions (SOAP)[28, 14, 41, 31, 32, 18, 21], Faber–Christensen–Huang–Lilienfeld (FCHL)[29, 30, 33, 34, 38], and similar descriptors offer atom-level descriptions within molecular or material environments based on physical insights and are effective in regressing chemical quantities, such as interatomic potentials (IAP) and nuclear magnetic resonance (NMR) chemical shifts.[28, 14, 41, 42, 29, 30, 15, 31, 32, 33, 34, 35, 36, 11, 37, 21, 38, 23, 24, 43] Notably, IAPs built using descriptors and Gaussian Process Regression (GPR) [14] have been termed Gaussian approximation potentials (GAP) and have found success in the exploration of the chemical space of molecules and materials.[14, 37, 21] Both kernel ridge regression (KRR) and GPR have been employed to improve the accuracy of NMR chemical shift prediction [44, 41, 42, 29, 30, 43]. However, the dimensionality of the descriptors becomes a barrier to generalization and high accuracy as the molecular or material composition becomes more diverse owing to the addition of different types of elements.[35, 19, 45, 39]
Recently, deep-learning models based on graph neural networks (GNNs) have been proposed to describe chemical spaces using graph representations.[46, 47, 48, 49, 50, 9, 17, 18, 51, 52, 19, 53, 20, 45, 54, 55, 56, 24, 25, 57] In most GNN-based IAPs, atoms within a molecular or material environment are represented as nodes, and their local connectivity as edges in a graph. The graph is then convolved to embed atom-specific information within each node, and further processed using multilayer perceptrons (MLP) to predict target observables. Numerous high-quality GNN-based IAPs have been developed for screening high-accuracy first-principles calculations and proposing molecular and crystal structures. Universal graph-based IAPs capable of comprehensively handling molecules and materials, such as MEGNet[52], M3GNet[10], Allegro[58], GNoME[59], and MACE[60, 61, 62] have emerged and are advancing rapidly.
Similarly, GNN-based models have been developed to predict NMR chemical shifts.[46, 47, 49, 50, 53, 57] For example, Yanfei Guan et al. achieved an accuracy of for H and for C by training the SchNet architecture [51] on molecular NMR chemical shifts (CASCADE)[46]. However, the scalability remains an issue due to the increasing optimization costs of GNN and MLP parameters when the size of datasets increase. Han et al. addressed this issue by constraining the nodes in a GNN to heavy elements only, thereby rendering the construction of scalable GNN-based NMR chemical shift models feasible while achieving a state-of-the-art prediction accuracy comparable to that of CASCADE.[53]
Notably, both descriptor-based and GNN methods face challenges. The former faces increased learning costs as the composition becomes more complex, and the latter faces increasing parameter optimization costs with larger training datasets. To address these issues simultaneously, we focused on the potential utility of the outputs from pre-trained GNN-based IAPs as descriptors. We considered these outputs GNN transfer learning (GNN-TF) descriptors and built machine-learning models for predicting chemical properties. Note that there are existing studies attempting to apply pre-trained GNN potentials to other tasks, particularly to generative modeling. [63, 64, 65, 66]
The remainder of this paper is organized as follows. Section II details the GNN-TF descriptor and the kernel method, implemented on both classical and quantum computers, for predicting NMR chemical shifts of H, C, N, O, and F. Section III presents the performance of our developed machine learning models. Section IV discusses the benefits and applications of the GNN-TF descriptor. Finally, Section V concludes the paper.
II Method:
Transfer Learning Using Pre-trained Graph Neural Network
In this section, we discuss the transfer learning of a pre-trained GNN-based IAP. This approach integrates the outputs from the GNN layer of the IAP as shown in Fig. 1.
The architecture of a GNN-based IAP can be broadly segmented into a GNN layer and an MLP layer (gray area of Fig. 1). For the IAP, we opted for two backbones: a MEGNet pre-trained on the QM9 dataset[67] and a M3GNet trained on the MPF.2021.2.8 dataset, which encompasses compounds covering all 89 elements from the Materials Project.[10]
The parameters of the GNN layer in the M3GNet IAP have been optimized for the prediction of system energy, forces, and stress tensors. When fed with the coordinates, s, of a molecule with atoms, the GNN layer generates a vector, , which mirrors the environment of the th atom in the molecule—henceforth referred to as the GNN-TF descriptor.
The GNN layer for both MEGNet and M3GNet outputs GNN-TF descriptors with dimensions of 32 and 64 per atom, respectively. Using GNN-TF descriptors as input, a regression model was constructed to predict NMR chemical shielding constants.
For the regressor, one can choose methodologies, such as GPR, KRR, or feed-forward neural network (NNs), which are contingent on the specific task. In this study, to ensure a maximally fair comparison with other descriptor-based techniques, we adopted KRR.
KRR combines the merits of ridge regression, which offers regularization to mitigate overfitting, with the kernel method, facilitating nonlinear regression. In kernel methods, the data —in the context of our study, the GNN-TF descriptors— are mapped into a high-dimensional feature space through a non-linear kernel function. The Laplacian kernel, which is known to produce favorable results for chemical properties prediction [68, 13, 29], was employed
where represents the hyperparameter of the kernel. In KRR, the predicted value for the target chemical property of the target atom is derived from the GNN-TF descriptor as follows:
Here, represents the element of the regression coefficient vector, , of size . The regression coefficients are determined by solving a ridge-regularized least-squares problem, which can be reduced to:
where denotes the identity matrix, denotes the chemical properties of each training data samples, and denotes the regularization parameter. The matrix , is a kernel matrix, with elements given by .
All computations related to the KRR were executed using Scikit-learn v.1.2.2,[69] and the hyperparameters of each model were tuned using Optuna v.2.10.[70] The hyperparameters of all models were obtained using only one random 80:20 split of the dataset. We employed five-fold cross-validation for all hyperparameter optimizations. For dataset sizes of up to 50K items, we conducted hyperparameter optimization for 100 iterations, whereas for datasets containing 100K items, we limited the optimization to 50 iterations.
The quantum-kernel method leverages quantum computers to compute kernels, [16, 71, 72] which is achieved by embedding feature vectors generated by classical computers into quantum states. This method calculates the inner product of these quantum states to derive the desired kernels. Embedding feature vectors into quantum states corresponds to mapping them onto a Hilbert space with dimensions raised to the power of two quantum bits (qubits). Using the kernel matrix constructed on a quantum computer, we performed a quantum kernel ridge regression (KRR), commonly referred to as quantum KRR (QKRR).
In this study, we adopted the natural parameterized quantum circuits (NPQC) Kernel, which has been demonstrated to possess performance characteristics similar to the Gaussian kernel, both theoretically and in actual hardware experiments [73, 74, 75]. All computations were conducted using Scikit-qulacs [69, 76, 77]. The quantum kernel was constructed in a 10-qubit space. Hyperparameters for the quantum kernel were determined through grid search. The determined parameters of NPQC kernel were = 1.5 and the repetition times of embedding 40. The regularization hyperparameter in QKRR was determined using 10 iterations of randomized search.
III Results
In Section III.1, because we deal with many elements, we compared the dimensional efficiency of our proposed GNN-TF descriptor to well-established physics-inspired descriptors. Note that the GNN-TF descriptor can better handle complex chemical systems by exploiting the GNN-based IAP architecture.
In Section III.2, we focused on the accuracy of the GNN-TF descriptor in predicting NMR chemical shifts, which are key to understanding molecular details (e.g., interatomic distances and bond angles). This scenario provides an ideal test for determining how well the GNN-TF descriptor works in our study.
Our analysis began by comparing quantum kernel learning, in which the kernels are tested using a quantum computer emulator with traditional kernel learning methods. We then checked the accuracy of the GNN-TF descriptors across the different pretrained GNN models.
Finally, we juxtaposed our GNN-TF descriptor using well-established physics-inspired descriptors. This comparison demonstrates the superiority of the proposed descriptor in terms of efficiency and accuracy. Furthermore, it highlights its potential for accurately predicting chemical properties, which is crucial for advancing research in the molecular and material sciences.
III.1 Dimensional Efficiency
| SOAP111SOAP and FCHL were generated by Dscribe 0.4.0[28] and QML 0.4.0.12[78], respectively. The default hyperparameters were selected as in QM9NMR paper. | FCHL1911footnotemark: 1 |
| Scaling with respect to | |||||
| 5740 | 740 | 64 | |||
| 22680 | 2440 | 64 | |||
| 1737120 | 162336 | 64 |
At the atomic level, descriptors are tools designed to encode information about atoms within molecules or crystalline materials into vectors. Popular descriptors, such as SOAP and FCHL18, excel at intricately capturing the environment within an atom’s cutoff radius.
Although these descriptors have achieved significant success in various accuracy benchmarks, they also present challenges due to their large dimensions. Various strategies have been developed to address these challenges,[79, 80, 81, 34] including refining the descriptor itself, using principal component analysis for dimensionality reduction, and exploring NNs to encode them. In particular, Christensen et al. applied Behler’s method of the atom-centered symmetry function [82] for NN potential to discretize FCHL18[33] to derive a compact and accurate FCHL19.[34]
In Table I, we present the scaling of the SOAP, FCHL19, and M3GNet GNN-TF descriptors in response to an increase in the number of elemental species considered. Additionally, for the QM9, QMugs,[83] and MPF.2021.8 datasets,[10] the descriptor dimensions corresponding to 5, 10, and 89 elemental species comprising each dataset are summarized, respectively.
Remarkably, with an increase in the number of element types, both SOAP and FCHL19 exhibited quadratic scaling. As a snapshot, when representing five elements in the QM9 dataset, the SOAP and FCHL19 methods have dimensions of 5,740 and 740, respectively. This dimensional disparity increases with the number of elemental types. Hence, to represent the 89 elements, the dimensions increased to 1,737,120 and 162,336, respectively. This is tens-of-thousands times larger than the concise 64 dimensions of the M3GNet GNN-TF descriptor.
Owing to its consistent dimensionality, irrespective of the increase in elements, the GNN-TF descriptor is overwhelmingly compact.
III.2 Prediction Accuracy: NMR Chemical Shifts
The NMR chemical shifts, , were predicted using the chemical shielding constant of the reference substance, , as the baseline. The NMR chemical shift was calculated using the following equation:
[…]
In our study, we utilized the QM9NMR dataset, which contains approximately 134K small organic molecules containing C, N, O, and F (excluding H), with each molecule having no more than nine atoms.[29, 67] This dataset provides the detailed NMR chemical shielding constants for these molecules. To analyze how the model accuracy changes with training data size, we adopted an approach similar to that used in the original publication of the QM9NMR dataset.[29] Specifically, for C, of a total of 831K data points, we randomly withheld 50K data points to build our test set. Subsequently, from the remaining C NMR chemical shifts, we randomly selected subsets containing 100, 200, 500, 1K, 2K, 5K, 10K, 50K, 100K, and 200K data points to create various training sets. For the other isotopes (i.e., H, N, O, and F), the test sets were similarly established by withholding 50K, 30K, 50K, and 1K data points, respectively. The training size for F was set to 2K, whereas the other isotopes were trained on datasets of 100K data points. In addition to the QM9 NMR dataset, we sought to validate the performance of our model on external datasets. Hence, we employed the two sets of molecules provided in another study;[29] one consisting of 40 drug molecules from the GDB17 universe and another containing 12 drugs with 17 or more heavy atoms.
Figure 2 shows the relationship between the mean absolute error (MAE) for C NMR shielding constant predictions and the training data size. Both QKRR and KRR demonstrated consistent improvements in predictive accuracy with an increase in training size. Notably, the quantum kernel exhibited a performance comparable to that of the Laplacian kernel.
For a training size of 100K, the MAE for the 13C predictions was . In a comparative study by Gupta et al., the KRR models using the Coulomb matrix (CM),[89] SOAP, and FCHL descriptors reported MAEs of approximately 4, 2.1, and , respectively, for the same training size.[29]
Compared with the CM descriptor, our GNN-TF descriptor showed significantly better predictive capabilities, achieving an MAE that was nearly half that of the CM descriptor. Although our method did not exceed the accuracy levels of SOAP and FCHL, the performance of the GNN-TF descriptor was competitive, highlighting its potential as a robust descriptor.
| GNN-TF descriptor/Regressor | RMSE (ppm) |
| SchNet/NN222 The value is taken from [90] | 12.8 |
| MEGNet/KRR | 20.4 |
| M3GNet/KRR | 10.0 |
Next, we compared the performance of the GNN-TF descriptors derived from different IAP architectures.
Recently, independent of our work, a predictive model for C NMR chemical shielding was proposed using a pretrained IAP known as SchNet, which is a pioneering GNN used as a descriptor.[90] This model was trained on 400 data points of C NMR chemical shielding constants of the molecules in QM9 dataset,[67] with the SchNet GNN-TF descriptor as an input to a feed-forward NN for regression. The predictive accuracy of the SchNet/NN was .
In pursuit of a fair comparison with their model, we applied Kernel Ridge Regression (KRR) using pre-trained M3GNet and MEGNet GNN-TF descriptors, setting our training data size to 400 data points of C NMR chemical shielding constants. The results of this comparative study are summarized in Table II.
The highest predictive accuracy was achieved by the M3GNet/KRR model, with an root mean-squared error of . This value surpassed the achieved by the MEGNet/KRR model and the achieved by the SchNet/NN model. This clearly underscores the superior performance of the M3GNet GNN descriptor over the other GNN-TF descriptors.
These results suggest that the architecture of the GNN-based IAP influences the performance of the GNN-TF descriptor. A more detailed discussion of the nuances of these architectural differences is presented in Section IV.1.
The accuracy of KRR models incorporating the GNN-TF descriptor for NMR chemical shifts was evaluated for molecules containing five elements that were not present in the training dataset. Table III lists the statistical performance metrics for predicting NMR chemical shifts. Across all elements, the MAE for the test set remained below . The MAE for H and F were notably low at and , respectively, indicating a high degree of prediction accuracy for these nuclei in the unseen molecules. The MAE for O, although higher at , still reflects a reasonable predictive capability, given the complexity of the oxygen chemical shifts. The standard deviation (STD) and interquartile range (IQR) values in the table represent the distribution of chemical shifts within the training data, rather than the accuracy of the model itself. Thus, the higher STD and IQR values for O do not indicate a lack of model precision but rather the natural variability inherent in the O chemical shifts within the training data. The MAE/STD ratio can still offer insights into model performance relative to data variability. For example, the relatively low ratio of O (2.21%) suggests that the model predictions are consistent with the diversity of the training data. Similarly, the higher ratios for H and F indicate that the accuracy of the model is commendable, particularly when considering the broad range of chemical shifts represented in the training dataset.
Subsequently, these models were employed to predict the NMR chemical shifts of molecules containing five elements that were not included in the training data. The results are shown in Fig. 3. The MAE for each nucleus were found to be for H, for C, for N, for O, and for F. The remarkably low MAE for H and C underscores the high accuracy of our model for these nuclei, with predictions that closely mirror the calculated values. The model performed well for the more challenging N and O nuclei, where the chemical shifts can be significantly affected by subtle changes in the molecular structure and environment, as indicated by the MAE values. The F nucleus, while having a higher MAE, showed excellent agreement with the DFT/GIAO calculations, suggesting that the model predictions were robust, even for nuclei with typically higher chemical shift ranges. These results demonstrate the strong predictive power and potential of the model as a reliable tool for accurately predicting NMR chemical shifts across a variety of nuclei, even in molecules beyond the scope of the training data.
| H | C | N | O | F | |
| MAE (ppm) | 0.18 | 2.28 | 3.42 | 4.95 | 2.65 |
| STD (ppm) | 1.98 | 51.96 | 119.58 | 224.40 | 34.07 |
| IQR (ppm) | 2.34 | 59.93 | 211.19 | 354.25 | 36.77 |
| MAE/STD () | 8.93 | 4.38 | 2.86 | 2.21 | 7.78 |
| FCHL333FCHL results are taken from [29]. | M3GNet GNN-TF | |
| 50K QM9 | 1.88 | 2.28 |
| 40 drugs | 3.7 | 3.47 |
| 12 drugs | 4.2 | 4.23 |
We then expanded our assessment to evaluate the predictive ability of our model for molecules larger than those in the QM9 NMR dataset. As such, we incorporate the test sets provided in another study,[29] which comprised 40 drug molecules from the GDB17 universe and another set containing 12 drugs with 17 or more heavy atoms.
Table IV presents the benchmark results for these test sets using our M3GNet GNN-TF descriptor. For comparison, we used the FCHL descriptor from Gupta’s study.[29] To ensure a fair comparison, we employed our M3GNet GNN-TF descriptor model trained on a size of 100K C. For both models, an increased molecular size in the dataset correlated with deterioration of the MAE value. Notably, although our M3GNet GNN-TF descriptor did not match the value achieved by the FCHL descriptor for the QM9 50K test set, our model exhibited an MAE value that was approximately lower for the 40 GDB17 dataset test. The results were nearly identical for the set of 12 drugs with 17 or more heavy atoms, highlighting that the GNN-TF descriptor was less affected by increasing molecular size.
For a detailed comparison, Fig. 4 illustrates the molecule-specific MAE values for both drug test sets. The molecular structures are provided in Appendix B, and more intricate details are available in another study.[29] Our M3GNet GNN-TF descriptor-based prediction model ensured that the highest MAE values for individual molecules across both test sets remained below . Intriguingly, the desflurane molecule, which posed the greatest challenge, showed MAE values of and for the FCHL and M3GNet GNN-TF descriptor models, respectively. This suggests an approximately 80% reduction in the MAE with our descriptor, which is likely attributable to differences in the encompassed descriptor domain.
[…]
In summary, the proposed M3GNet GNN-TF descriptor demonstrates the capability of predicting C NMR chemical shifts for molecules outside the training dataset with an accuracy comparable to that of the state-of-the-art FCHL descriptor.
IV Discussion
IV.1 Influence of Architectural Choices on GNN-TF Descriptor Performance
In our exploration of different architectures for generating GNN-TF descriptors, we observed several patterns. SchNet, which operates on GNN-based local descriptors to evaluate systems as summations of atomic energies, accounts only for pairwise interactions. This limited inclusion could potentially constrain expressions, leading to inadequate representational power.
The subpar performance of MEGNet during transfer learning may be attributed to its architectural design as it integrates atomic (local) descriptors into molecular (global) descriptors through concatenation. This means that the final piece of information passed to the MLP is not extracted directly from the end of the model, which might not be the optimal representation for targeted atomic-wise property prediction; however, it is expected to be suitable for molecule-wise property predictions.
IV.2 Significance of Dataset Size and Diversity
The M3GNet training regimen incorporates data from 187,687 ionic steps spanning 62,783 compounds, including 187,687 energies, 16,875,138 force components, and 1,689,183 stress components. This diverse dataset covers 89 elements from the periodic table. The model is not limited to learning only the energies associated with these elements but extends to atomic-level forces. Moreover, M3GNet training includes not only stable structures but also the processes of structure optimization. The ingestion of vast amounts of data from crystalline systems may have endowed the M3GNet with enhanced expression, potentially making it adept at interpolating molecular systems.
IV.3 Potential for Transfer Learning on Quantum Computer
There is a potential for leveraging quantum computation approaches.[91] Specifically, our 10 qubit QKRR, facilitated by a simulator, demonstrated a performance comparable to that of state-of-the-art KRR. This is underpinned by the theoretical equivalence of the NPQC with the RBF.
The quantum kernel method stands out because of its capability to compute with fewer measurement iterations than other quantum computation methodologies, such as quantum neural networks.[92] In particular, our proposed M3GNet GNN-TF descriptor can be feasibly realized with a minimum of six qubits, enabling evaluations with a quantum bit count that is more efficient than traditional descriptors, such as SOAP. However, embedding for higher-dimensional SOAP appears to be a challenge, possibly due to noise.
From a futuristic perspective, there is excitement about the possibility of developing kernels that traditional computers cannot express, as well as accelerating the inversion calculations of kernel matrices using quantum algorithms. The constant scaling property of our proposed method concerning element number dimensions may significantly contribute to real-time material exploration powered by quantum-classical hybrid algorithms in the near future.
V Conclusion
The dynamics of machine learning and its extensive applications across various domains are driving cutting-edge research. Our endeavor to integrate transfer learning with pretrained IAP GNNs for NMR chemical shift prediction offers a paradigm shift in efficiency and scalability. The GNN-TF descriptor presents an unparalleled advantage in terms of scalability due to its consistent dimensionality, irrespective of the number of elements.
Comparative evaluations with other renowned descriptors, such as SOAP, suggest that the GNN-TF descriptor can match, if not surpass, the performance of its contemporaries while maintaining a more compact representation. This is especially important when factoring large datasets, where dimensionality can exponentially burgeon.
Architectural choice plays a pivotal role in the performance of GNN-TF descriptors. Moreover, the diversity and vastness of the training dataset, which encompasses myriad elemental types and structural configurations, augment the robustness and versatility of the GNN.
Our proposed model has immense potential for creating a unified framework capable of predicting various atomic and molecular properties simultaneously, presenting profound implications for accelerated material and molecular research. This potential union of multiple predictions can usher in an era of comprehensive understanding and quicker innovations, possibly revolutionizing fields, such as catalysis, drug discovery, and material design.
The union of transfer learning with pretrained GNNs not only augments prediction accuracy but also drastically reduces learning costs, presenting a cost-effective and efficient alternative to more computationally intensive methods. As we move toward an era in which data-driven insights and models govern the pace of innovation, our research offers a promising pathway for future endeavors in the domain of chemical property predictions with both classical and quantum computers.
Note added - As we were finalizing this manuscript, we became aware of recent articles [90, 93, 94] that also utilize intermediate information from graph neural network potentials. In Section III.2, we added a direct comparison between our results and theirs. Elijošius et al. applied the pre-trained MACE descriptor to generative modeling of molecules [94].
Acknowledgements.
We thank Nobuki Inoue and Tuan Minh Do for fruitful discussions. This project was supported by funding from the MEXT Quantum Leap Flagship Program (MEXTQLEAP) through Grant No. JPMXS0120319794, and the JST COI-NEXT Program through Grant No. JPMJPF2014. The completion of this research was partially facilitated by the JSPS Grants-in-Aid for Scientific Research (KAKENHI), specifically Grant Nos. JP23H03819 and JP21K18933. We thank the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo for the use of the facilities. This work was also achieved through the use of SQUID at the Cybermedia Center, Osaka University.References
- Reymond et al. [2010] J.-L. Reymond, R. van Deursen, L. C. Blum, and L. Ruddigkeit, Chemical space as a source for new drugs, MedChemComm 1, 30 (2010).
- Ruddigkeit et al. [2012] L. Ruddigkeit, R. van Deursen, L. C. Blum, and J.-L. Reymond, Enumeration of 166 billion organic small molecules in the chemical universe database GDB-17, Journal of Chemical Information and Modeling 52, 2864 (2012).
- Gómez-Bombarelli et al. [2016] R. Gómez-Bombarelli, J. Aguilera-Iparraguirre, T. D. Hirzel, D. Duvenaud, D. Maclaurin, M. A. Blood-Forsythe, H. S. Chae, M. Einzinger, D.-G. Ha, T. Wu, G. Markopoulos, S. Jeon, H. Kang, H. Miyazaki, M. Numata, S. Kim, W. Huang, S. I. Hong, M. Baldo, R. P. Adams, and A. Aspuru-Guzik, Design of efficient molecular organic light-emitting diodes by a high-throughput virtual screening and experimental approach, Nature Materials 15, 1120 (2016).
- Curtarolo et al. [2013] S. Curtarolo, G. L. W. Hart, M. B. Nardelli, N. Mingo, S. Sanvito, and O. Levy, The high-throughput highway to computational materials design, Nature Materials 12, 191 (2013).
- Nie et al. [2023] W. Nie, Q. Wan, J. Sun, M. Chen, M. Gao, and S. Chen, Ultra-high-throughput mapping of the chemical space of asymmetric catalysis enables accelerated reaction discovery, Nature Communications 14, 10.1038/s41467-023-42446-5 (2023).
- Oganov et al. [2019] A. R. Oganov, C. J. Pickard, Q. Zhu, and R. J. Needs, Structure prediction drives materials discovery, Nature Reviews Materials 4, 331 (2019).
- Pierens [2014] G. K. Pierens, 1h and 13c nmr scaling factors for the calculation of chemical shifts in commonly used solvents using density functional theory, Journal of computational chemistry 35, 1388 (2014).
- Hartman et al. [2016] J. D. Hartman, R. A. Kudla, G. M. Day, L. J. Mueller, and G. J. O. Beran, Benchmark fragment-based 1h, 13c, 15n and 17o chemical shift predictions in molecular crystals, Physical Chemistry Chemical Physics 18, 21686 (2016).
- Büning and Grimme [2023] J. B. K. Büning and S. Grimme, Computation of CCSD(t)-quality NMR chemical shifts via machine learning from DFT, Journal of Chemical Theory and Computation 19, 3601 (2023).
- Chen and Ong [2022] C. Chen and S. P. Ong, A universal graph deep learning interatomic potential for the periodic table, Nature Computational Science 2, 718 (2022).
- Lodewyk et al. [2011] M. W. Lodewyk, M. R. Siebert, and D. J. Tantillo, Computational prediction of 1h and 13c chemical shifts: A useful tool for natural product, mechanistic, and synthetic organic chemistry, Chemical Reviews 112, 1839 (2011).
- Lauro et al. [2020] G. Lauro, P. Das, R. Riccio, D. S. Reddy, and G. Bifulco, Dft/nmr approach for the configuration assignment of groups of stereoisomers by the combination and comparison of experimental and predicted sets of data, The Journal of Organic Chemistry 85, 3297–3306 (2020).
- Hansen et al. [2015] K. Hansen, F. Biegler, R. Ramakrishnan, W. Pronobis, O. A. Von Lilienfeld, K.-R. Muller, and A. Tkatchenko, Machine learning predictions of molecular properties: Accurate many-body potentials and nonlocality in chemical space, The journal of physical chemistry letters 6, 2326 (2015).
- Deringer et al. [2021] V. L. Deringer, A. P. Bartók, N. Bernstein, D. M. Wilkins, M. Ceriotti, and G. Csányi, Gaussian process regression for materials and molecules, Chemical Reviews 121, 10073 (2021).
- Faber et al. [2017] F. A. Faber, L. Hutchison, B. Huang, J. Gilmer, S. S. Schoenholz, G. E. Dahl, O. Vinyals, S. Kearnes, P. F. Riley, and O. A. von Lilienfeld, Prediction errors of molecular machine learning models lower than hybrid dft error, Journal of Chemical Theory and Computation 13, 5255 (2017).
- Sajjan et al. [2022] M. Sajjan, J. Li, R. Selvarajan, S. H. Sureshbabu, S. S. Kale, R. Gupta, V. Singh, and S. Kais, Quantum machine learning for chemistry and physics, Chemical Society Reviews 51, 6475 (2022).
- Keith et al. [2021] J. A. Keith, V. Vassilev-Galindo, B. Cheng, S. Chmiela, M. Gastegger, K.-R. Müller, and A. Tkatchenko, Combining machine learning and computational chemistry for predictive insights into chemical systems, Chemical Reviews 121, 9816 (2021).
- Wan et al. [2023] K. Wan, J. He, and X. Shi, Construction of high accuracy machine learning interatomic potential for surface/interface of nanomaterials—a review, Advanced Materials 10.1002/adma.202305758 (2023).
- Wu et al. [2018] Z. Wu, B. Ramsundar, E. N. Feinberg, J. Gomes, C. Geniesse, A. S. Pappu, K. Leswing, and V. Pande, MoleculeNet: a benchmark for molecular machine learning, Chemical Science 9, 513 (2018).
- Reiser et al. [2022] P. Reiser, M. Neubert, A. Eberhard, L. Torresi, C. Zhou, C. Shao, H. Metni, C. van Hoesel, H. Schopmans, T. Sommer, and P. Friederich, Graph neural networks for materials science and chemistry, Communications Materials 3, 10.1038/s43246-022-00315-6 (2022).
- Bartók et al. [2017] A. P. Bartók, S. De, C. Poelking, N. Bernstein, J. R. Kermode, G. Csányi, and M. Ceriotti, Machine learning unifies the modeling of materials and molecules, Science Advances 3, 10.1126/sciadv.1701816 (2017).
- Kocer et al. [2022] E. Kocer, T. W. Ko, and J. Behler, Neural network potentials: A concise overview of methods, Annual Review of Physical Chemistry 73, 163–186 (2022).
- Langer et al. [2022] M. F. Langer, A. Goeßmann, and M. Rupp, Representations of molecules and materials for interpolation of quantum-mechanical simulations via machine learning, npj Computational Materials 8, 41 (2022).
- Liu et al. [2021] Z. Liu, L. Lin, Q. Jia, Z. Cheng, Y. Jiang, Y. Guo, and J. Ma, Transferable multilevel attention neural network for accurate prediction of quantum chemistry properties via multitask learning, Journal of Chemical Information and Modeling 61, 1066–1082 (2021).
- Merchant et al. [2023a] A. Merchant, S. Batzner, S. S. Schoenholz, M. Aykol, G. Cheon, and E. D. Cubuk, Scaling deep learning for materials discovery, Nature 624, 80–85 (2023a).
- Gao et al. [2020] P. Gao, J. Zhang, Q. Peng, J. Zhang, and V.-A. Glezakou, General protocol for the accurate prediction of molecular 13c/1h nmr chemical shifts via machine learning augmented dft, Journal of Chemical Information and Modeling 60, 3746 (2020).
- Macalino et al. [2015] S. J. Y. Macalino, V. Gosu, S. Hong, and S. Choi, Role of computer-aided drug design in modern drug discovery, Archives of Pharmacal Research 38, 1686 (2015).
- Himanen et al. [2020] L. Himanen, M. O. Jäger, E. V. Morooka, F. Federici Canova, Y. S. Ranawat, D. Z. Gao, P. Rinke, and A. S. Foster, Dscribe: Library of descriptors for machine learning in materials science, Computer Physics Communications 247, 106949 (2020).
- Gupta et al. [2021] A. Gupta, S. Chakraborty, and R. Ramakrishnan, Revving up 13c NMR shielding predictions across chemical space: benchmarks for atoms-in-molecules kernel machine learning with new data for 134 kilo molecules, Machine Learning: Science and Technology 2, 035010 (2021).
- Gerrard et al. [2020] W. Gerrard, L. A. Bratholm, M. J. Packer, A. J. Mulholland, D. R. Glowacki, and C. P. Butts, IMPRESSION – prediction of NMR parameters for 3-dimensional chemical structures using machine learning with near quantum chemical accuracy, Chemical Science 11, 508 (2020).
- Willatt et al. [2019] M. J. Willatt, F. Musil, and M. Ceriotti, Atom-density representations for machine learning, The Journal of Chemical Physics 150, 10.1063/1.5090481 (2019).
- Bartók et al. [2013] A. P. Bartók, R. Kondor, and G. Csányi, On representing chemical environments, Physical Review B 87, 10.1103/physrevb.87.184115 (2013).
- Faber et al. [2018] F. A. Faber, A. S. Christensen, B. Huang, and O. A. von Lilienfeld, Alchemical and structural distribution based representation for universal quantum machine learning, The Journal of Chemical Physics 148, 10.1063/1.5020710 (2018).
- Christensen et al. [2020] A. S. Christensen, L. A. Bratholm, F. A. Faber, and O. A. von Lilienfeld, FCHL revisited: Faster and more accurate quantum machine learning, The Journal of Chemical Physics 152, 10.1063/1.5126701 (2020).
- Kabylda et al. [2023] A. Kabylda, V. Vassilev-Galindo, S. Chmiela, I. Poltavsky, and A. Tkatchenko, Efficient interatomic descriptors for accurate machine learning force fields of extended molecules, Nature Communications 14, 10.1038/s41467-023-39214-w (2023).
- Musil et al. [2021] F. Musil, A. Grisafi, A. P. Bartók, C. Ortner, G. Csányi, and M. Ceriotti, Physics-inspired structural representations for molecules and materials, Chemical Reviews 121, 9759 (2021).
- Bartók et al. [2010] A. P. Bartók, M. C. Payne, R. Kondor, and G. Csányi, Gaussian approximation potentials: The accuracy of quantum mechanics, without the electrons, Physical Review Letters 104, 10.1103/physrevlett.104.136403 (2010).
- Parsaeifard et al. [2021] B. Parsaeifard, D. Sankar De, A. S. Christensen, F. A. Faber, E. Kocer, S. De, J. Behler, O. Anatole von Lilienfeld, and S. Goedecker, An assessment of the structural resolution of various fingerprints commonly used in machine learning, Machine Learning: Science and Technology 2, 015018 (2021).
- Khan et al. [2023] D. Khan, S. Heinen, and O. A. von Lilienfeld, Kernel based quantum machine learning at record rate: Many-body distribution functionals as compact representations, The Journal of Chemical Physics 159, 10.1063/5.0152215 (2023).
- Rupp et al. [2015] M. Rupp, R. Ramakrishnan, and O. A. von Lilienfeld, Machine learning for quantum mechanical properties of atoms in molecules, The Journal of Physical Chemistry Letters 6, 3309 (2015).
- Cordova et al. [2022] M. Cordova, E. A. Engel, A. Stefaniuk, F. Paruzzo, A. Hofstetter, M. Ceriotti, and L. Emsley, A machine learning model of chemical shifts for chemically and structurally diverse molecular solids, The Journal of Physical Chemistry C 126, 16710 (2022).
- Kohlhoff et al. [2009] K. J. Kohlhoff, P. Robustelli, A. Cavalli, X. Salvatella, and M. Vendruscolo, Fast and accurate predictions of protein NMR chemical shifts from interatomic distances, Journal of the American Chemical Society 131, 13894 (2009).
- Tsitsvero et al. [2023] M. Tsitsvero, J. Pirillo, Y. Hijikata, and T. Komatsuzaki, Nmr spectrum prediction for dynamic molecules by machine learning: A case study of trefoil knot molecule, The Journal of Chemical Physics 158, 10.1063/5.0147398 (2023).
- Paruzzo et al. [2018] F. M. Paruzzo, A. Hofstetter, F. Musil, S. De, M. Ceriotti, and L. Emsley, Chemical shifts in molecular solids by machine learning, Nature Communications 9, 10.1038/s41467-018-06972-x (2018).
- Fung et al. [2021] V. Fung, J. Zhang, E. Juarez, and B. G. Sumpter, Benchmarking graph neural networks for materials chemistry, npj Computational Materials 7, 10.1038/s41524-021-00554-0 (2021).
- Guan et al. [2021] Y. Guan, S. V. S. Sowndarya, L. C. Gallegos, P. C. S. John, and R. S. Paton, Real-time prediction of 1h and 13c chemical shifts with DFT accuracy using a 3d graph neural network, Chemical Science 12, 12012 (2021).
- Liu et al. [2019] S. Liu, J. Li, K. C. Bennett, B. Ganoe, T. Stauch, M. Head-Gordon, A. Hexemer, D. Ushizima, and T. Head-Gordon, Multiresolution 3d-DenseNet for chemical shift prediction in NMR crystallography, The Journal of Physical Chemistry Letters 10, 4558 (2019).
- Han and Choi [2021] H. Han and S. Choi, Transfer learning from simulation to experimental data: NMR chemical shift predictions, The Journal of Physical Chemistry Letters 12, 3662 (2021).
- Kwon et al. [2020] Y. Kwon, D. Lee, Y.-S. Choi, M. Kang, and S. Kang, Neural message passing for NMR chemical shift prediction, Journal of Chemical Information and Modeling 60, 2024 (2020).
- Jonas and Kuhn [2019] E. Jonas and S. Kuhn, Rapid prediction of NMR spectral properties with quantified uncertainty, Journal of Cheminformatics 11, 10.1186/s13321-019-0374-3 (2019).
- Schütt et al. [2017] K. T. Schütt, P.-J. Kindermans, H. E. Sauceda, S. Chmiela, A. Tkatchenko, and K.-R. Müller, Schnet: A continuous-filter convolutional neural network for modeling quantum interactions (2017), arXiv:1706.08566 [stat.ML] .
- Chen et al. [2019] C. Chen, W. Ye, Y. Zuo, C. Zheng, and S. P. Ong, Graph networks as a universal machine learning framework for molecules and crystals, Chemistry of Materials 31, 3564 (2019).
- Han et al. [2022] J. Han, H. Kang, S. Kang, Y. Kwon, D. Lee, and Y.-S. Choi, Scalable graph neural network for NMR chemical shift prediction, Physical Chemistry Chemical Physics 24, 26870 (2022).
- Xu et al. [2019] K. Xu, W. Hu, J. Leskovec, and S. Jegelka, How powerful are graph neural networks? (2019), arXiv:1810.00826 [cs.LG] .
- Stärk et al. [2022] H. Stärk, D. Beaini, G. Corso, P. Tossou, C. Dallago, S. Günnemann, and P. Lió, 3D infomax improves GNNs for molecular property prediction, in Proceedings of the 39th International Conference on Machine Learning, Proceedings of Machine Learning Research, Vol. 162, edited by K. Chaudhuri, S. Jegelka, L. Song, C. Szepesvari, G. Niu, and S. Sabato (PMLR, 2022) pp. 20479–20502.
- Jiang et al. [2021] D. Jiang, Z. Wu, C.-Y. Hsieh, G. Chen, B. Liao, Z. Wang, C. Shen, D. Cao, J. Wu, and T. Hou, Could graph neural networks learn better molecular representation for drug discovery? a comparison study of descriptor-based and graph-based models, Journal of Cheminformatics 13, 10.1186/s13321-020-00479-8 (2021).
- Kang et al. [2020] S. Kang, Y. Kwon, D. Lee, and Y.-S. Choi, Predictive modeling of nmr chemical shifts without using atomic-level annotations, Journal of Chemical Information and Modeling 60, 3765–3769 (2020).
- Musaelian et al. [2023] A. Musaelian, S. Batzner, A. Johansson, L. Sun, C. J. Owen, M. Kornbluth, and B. Kozinsky, Learning local equivariant representations for large-scale atomistic dynamics, Nature Communications 14, 579 (2023).
- Merchant et al. [2023b] A. Merchant, S. Batzner, S. S. Schoenholz, M. Aykol, G. Cheon, and E. D. Cubuk, Scaling deep learning for materials discovery, Nature 10.1038/s41586-023-06735-9 (2023b).
- Batatia et al. [2022] I. Batatia, D. P. Kovacs, G. N. C. Simm, C. Ortner, and G. Csanyi, MACE: Higher order equivariant message passing neural networks for fast and accurate force fields, in Advances in Neural Information Processing Systems, edited by A. H. Oh, A. Agarwal, D. Belgrave, and K. Cho (2022).
- Kovács et al. [2023] D. P. Kovács, J. H. Moore, N. J. Browning, I. Batatia, J. T. Horton, V. Kapil, I.-B. Magdău, D. J. Cole, and G. Csányi, Mace-off23: Transferable machine learning force fields for organic molecules, arXiv preprint arXiv:2312.15211 (2023).
- Batatia et al. [2023] I. Batatia, P. Benner, Y. Chiang, A. M. Elena, D. P. Kovács, J. Riebesell, X. R. Advincula, M. Asta, W. J. Baldwin, N. Bernstein, A. Bhowmik, S. M. Blau, V. Cărare, J. P. Darby, S. De, F. D. Pia, V. L. Deringer, R. Elijošius, Z. El-Machachi, E. Fako, A. C. Ferrari, A. Genreith-Schriever, J. George, R. E. A. Goodall, C. P. Grey, S. Han, W. Handley, H. H. Heenen, K. Hermansson, C. Holm, J. Jaafar, S. Hofmann, K. S. Jakob, H. Jung, V. Kapil, A. D. Kaplan, N. Karimitari, N. Kroupa, J. Kullgren, M. C. Kuner, D. Kuryla, G. Liepuoniute, J. T. Margraf, I.-B. Magdău, A. Michaelides, J. H. Moore, A. A. Naik, S. P. Niblett, S. W. Norwood, N. O’Neill, C. Ortner, K. A. Persson, K. Reuter, A. S. Rosen, L. L. Schaaf, C. Schran, E. Sivonxay, T. K. Stenczel, V. Svahn, C. Sutton, C. van der Oord, E. Varga-Umbrich, T. Vegge, M. Vondrák, Y. Wang, W. C. Witt, F. Zills, and G. Csányi, A foundation model for atomistic materials chemistry (2023), arXiv:2401.00096 [physics.chem-ph] .
- Xie et al. [2021] T. Xie, X. Fu, O.-E. Ganea, R. Barzilay, and T. Jaakkola, Crystal diffusion variational autoencoder for periodic material generation, arXiv preprint arXiv:2110.06197 (2021).
- Wu et al. [2022] L. Wu, C. Gong, X. Liu, M. Ye, and Q. Liu, Diffusion-based molecule generation with informative prior bridges, Advances in Neural Information Processing Systems 35, 36533 (2022).
- Zaidi et al. [2022] S. Zaidi, M. Schaarschmidt, J. Martens, H. Kim, Y. W. Teh, A. Sanchez-Gonzalez, P. Battaglia, R. Pascanu, and J. Godwin, Pre-training via denoising for molecular property prediction, arXiv preprint arXiv:2206.00133 (2022).
- Jia et al. [2024] S. Jia, A. R. Parthasarathy, R. Feng, G. Cong, C. Zhang, and V. Fung, Derivative-based pre-training of graph neural networks for materials property predictions, Digital Discovery (2024).
- Ramakrishnan et al. [2014] R. Ramakrishnan, P. O. Dral, M. Rupp, and O. A. von Lilienfeld, Quantum chemistry structures and properties of 134 kilo molecules, Scientific Data 1, 10.1038/sdata.2014.22 (2014).
- Dral et al. [2015] P. O. Dral, O. A. von Lilienfeld, and W. Thiel, Machine learning of parameters for accurate semiempirical quantum chemical calculations, Journal of chemical theory and computation 11, 2120 (2015).
- Pedregosa et al. [2011] F. Pedregosa, G. Varoquaux, A. Gramfort, V. Michel, B. Thirion, O. Grisel, M. Blondel, P. Prettenhofer, R. Weiss, V. Dubourg, J. Vanderplas, A. Passos, D. Cournapeau, M. Brucher, M. Perrot, and E. Duchesnay, Scikit-learn: Machine learning in Python, Journal of Machine Learning Research 12, 2825 (2011).
- Akiba et al. [2019] T. Akiba, S. Sano, T. Yanase, T. Ohta, and M. Koyama, Optuna: A next-generation hyperparameter optimization framework (2019), arXiv:1907.10902 [cs.LG] .
- Schuld and Killoran [2019] M. Schuld and N. Killoran, Quantum machine learning in feature hilbert spaces, Physical Review Letters 122, 10.1103/physrevlett.122.040504 (2019).
- Kusumoto et al. [2021] T. Kusumoto, K. Mitarai, K. Fujii, M. Kitagawa, and M. Negoro, Experimental quantum kernel trick with nuclear spins in a solid, npj Quantum Information 7, 10.1038/s41534-021-00423-0 (2021).
- Haug et al. [2023] T. Haug, C. N. Self, and M. S. Kim, Quantum machine learning of large datasets using randomized measurements, Machine Learning: Science and Technology 4, 015005 (2023).
- Haug and Kim [2022] T. Haug and M. S. Kim, Natural parametrized quantum circuit, Physical Review A 106, 10.1103/physreva.106.052611 (2022).
- Benedetti et al. [2019] M. Benedetti, E. Lloyd, S. Sack, and M. Fiorentini, Parameterized quantum circuits as machine learning models, Quantum Science and Technology 4, 043001 (2019).
- [76] https://github.com/Qulacs-Osaka/scikit-qulacs.
- Suzuki et al. [2021] Y. Suzuki, Y. Kawase, Y. Masumura, Y. Hiraga, M. Nakadai, J. Chen, K. M. Nakanishi, K. Mitarai, R. Imai, S. Tamiya, T. Yamamoto, T. Yan, T. Kawakubo, Y. O. Nakagawa, Y. Ibe, Y. Zhang, H. Yamashita, H. Yoshimura, A. Hayashi, and K. Fujii, Qulacs: a fast and versatile quantum circuit simulator for research purpose, Quantum 5, 559 (2021).
- Christensen et al. [2017] A. S. Christensen, L. A. Bratholm, F. A. Faber, B. Huang, A. Tkatchenko, K. R. Müller, and O. A. von Lilienfeld, Qml: A python toolkit for quantum machine learning, https://github.com/qmlcode/qml (2017).
- Lopanitsyna et al. [2023] N. Lopanitsyna, G. Fraux, M. A. Springer, S. De, and M. Ceriotti, Modeling high-entropy transition metal alloys with alchemical compression, Physical Review Materials 7, 045802 (2023).
- Willatt et al. [2018] M. J. Willatt, F. Musil, and M. Ceriotti, Feature optimization for atomistic machine learning yields a data-driven construction of the periodic table of the elements, Physical Chemistry Chemical Physics 20, 29661 (2018).
- Li et al. [2022] S. Li, Y. Liu, D. Chen, Y. Jiang, Z. Nie, and F. Pan, Encoding the atomic structure for machine learning in materials science, Wiley Interdisciplinary Reviews: Computational Molecular Science 12, e1558 (2022).
- Behler [2011] J. Behler, Atom-centered symmetry functions for constructing high-dimensional neural network potentials, The Journal of Chemical Physics 134, 074106 (2011).
- Isert et al. [2022] C. Isert, K. Atz, J. Jiménez-Luna, and G. Schneider, Qmugs, quantum mechanical properties of drug-like molecules, Scientific Data 9, 273 (2022).
- Xin et al. [2017] D. Xin, C. A. Sader, U. Fischer, K. Wagner, P.-J. Jones, M. Xing, K. R. Fandrick, and N. C. Gonnella, Systematic investigation of dft-giao 15n nmr chemical shift prediction using b3lyp/cc-pvdz: application to studies of regioisomers, tautomers, protonation states and n-oxides, Organic & Biomolecular Chemistry 15, 928–936 (2017).
- Puzzarini et al. [2009] C. Puzzarini, G. Cazzoli, M. E. Harding, J. Vázquez, and J. Gauss, A new experimental absolute nuclear magnetic shielding scale for oxygen based on the rotational hyperfine structure of h2o17, The Journal of Chemical Physics 131, 10.1063/1.3274062 (2009).
- Wasylishen and Bryce [2002] R. E. Wasylishen and D. L. Bryce, A revised experimental absolute magnetic shielding scale for oxygen, The Journal of Chemical Physics 117, 10061–10066 (2002).
- Rosenau et al. [2018] C. P. Rosenau, B. J. Jelier, A. D. Gossert, and A. Togni, Exposing the origins of irreproducibility in fluorine nmr spectroscopy, Angewandte Chemie International Edition 57, 9528 (2018).
- Frisch et al. [2016] M. J. Frisch, G. W. Trucks, H. B. Schlegel, G. E. Scuseria, M. A. Robb, J. R. Cheeseman, G. Scalmani, V. Barone, G. A. Petersson, H. Nakatsuji, X. Li, M. Caricato, A. V. Marenich, J. Bloino, B. G. Janesko, R. Gomperts, B. Mennucci, H. P. Hratchian, J. V. Ortiz, A. F. Izmaylov, J. L. Sonnenberg, D. Williams-Young, F. Ding, F. Lipparini, F. Egidi, J. Goings, B. Peng, A. Petrone, T. Henderson, D. Ranasinghe, V. G. Zakrzewski, J. Gao, N. Rega, G. Zheng, W. Liang, M. Hada, M. Ehara, K. Toyota, R. Fukuda, J. Hasegawa, M. Ishida, T. Nakajima, Y. Honda, O. Kitao, H. Nakai, T. Vreven, K. Throssell, J. A. Montgomery, Jr., J. E. Peralta, F. Ogliaro, M. J. Bearpark, J. J. Heyd, E. N. Brothers, K. N. Kudin, V. N. Staroverov, T. A. Keith, R. Kobayashi, J. Normand, K. Raghavachari, A. P. Rendell, J. C. Burant, S. S. Iyengar, J. Tomasi, M. Cossi, J. M. Millam, M. Klene, C. Adamo, R. Cammi, J. W. Ochterski, R. L. Martin, K. Morokuma, O. Farkas, J. B. Foresman, and D. J. Fox, Gaussian˜16 Revision C.01 (2016), gaussian Inc. Wallingford CT.
- Rupp et al. [2012] M. Rupp, A. Tkatchenko, K.-R. Müller, and O. A. von Lilienfeld, Fast and accurate modeling of molecular atomization energies with machine learning, Physical Review Letters 108, 10.1103/physrevlett.108.058301 (2012).
- Samman et al. [2023] A. M. E. Samman, S. D. Castro, B. Morton, and S. D. Baerdemacker, Transfer learning graph representations of molecules for pka, 13c-nmr, and solubility 10.26434/chemrxiv-2023-hfcm5 (2023).
- Cerezo et al. [2021] M. Cerezo, A. Arrasmith, R. Babbush, S. C. Benjamin, S. Endo, K. Fujii, J. R. McClean, K. Mitarai, X. Yuan, L. Cincio, and P. J. Coles, Variational quantum algorithms, Nature Reviews Physics 3, 625 (2021).
- Mitarai et al. [2018] K. Mitarai, M. Negoro, M. Kitagawa, and K. Fujii, Quantum circuit learning, Physical Review A 98, 10.1103/physreva.98.032309 (2018).
- El-Samman et al. [2023] A. El-Samman, I. Husain, M. Huynh, S. De Castro, B. Morton, G. Acke, and S. De Baerdemacker, Global interpretability and geometry of graph convolu- tional neural networks for chemistry in terms of chemical moieties 10.26434/chemrxiv-2023-zhlmr (2023).
- Elijošius et al. [2024] R. Elijošius, F. Zills, I. Batatia, S. W. Norwood, D. P. Kovács, C. Holm, and G. Csányi, Zero shot molecular generation via similarity kernels, arXiv preprint arXiv:2402.08708 (2024).
Appendix A Distribution of Datasets for Each NMR Chemical Shift Prediction Model
The distributions of the training and test sets sampled from the QM9NMR dataset are shown in Fig. 4. Figure 4.(b) shows that above 5K, the distribution is in good agreement with the overall distribution of the C NMR shielding constants. For the other elemental species, the distributions of the training and test sets were in good agreement with the overall distribution.
Appendix B Validation of Learning Accuracy of NMR Chemical Shift Prediction
Figure 5 illustrates the accuracy of the KRR models trained using the M3GNet GNN-TF descriptor for five elemental species. The MAE values for the NMR shielding constant, for train/test, are as follows: for , 0.0344/0.1767; , 0.1420/2.2798 ppm; , 0.3910/3.4157 ppm; 0.8881/4.9509 ppm; and 0.0864/2.6518 ppm.
The structures of two sets of drug molecules [29] used for validating the constructed model beyond the QM9NMR dataset are depicted in Figs. 6 and 7, respectively. The predicted C NMR shielding constants for each drug molecule using the M3GNet GNN-TF/KRR model are shown in Fig. 8. These predictions are accompanied by the values predicted by the FCHL/KRR model.[29]